# Optimization and Generalization Guarantees for Weight Normalization

### v1

@@ 5 Optimization guarantees for WeightNorm
AB-PCV-LZ-MB:22] Let $\alpha_{\theta_{t}},\beta_{\theta_{t}}$ be as in Lemmas 5.1 and 5.2 respectively, and $B_{t}:=Q^{\theta}_{\kappa}\cap B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})\cap\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}$. Let $\theta^{*}\in\operatorname{arginf}_{\theta\in\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}}{\cal L}(\theta)$, $\overline{\theta}_{t+1}\in\operatorname{arginf}_{\theta\in B_{t}}{\cal L}(\theta)$, and $\gamma_{t}:=\frac{{\cal L}(\overline{\theta}_{t+1})-{\cal L}(\theta^{*})}{{\cal L}(\theta_{t})-{\cal L}(\theta^{*})}$. Under Assumptions 12, and 3, we have $\gamma_{t}\in[0,1)$, and for gradient descent with step size $\eta_{t}=\frac{\omega_{t}}{\beta_{\theta_{t}}},\omega_{t}\in(0,2)$,
${\cal L}(\theta_{t+1})-{\cal L}(\theta^{*})\leq\left(1-\frac{\alpha_{\theta_{t}}}{\beta_{\theta_{t}}}\omega_{t}(1-\gamma_{t})(2-\omega_{t})\right)({\cal L}(\theta_{t})-{\cal L}(\theta^{*}))~{}.$ (18)
The result in (18) implies training convergence as the number of iterations increase. Since the rate at which the loss decreases is time- and weight-dependent in (18), it is difficult to establish the number of steps needed for convergence to a specific loss value; however, this shows that our framework may cover cases where the convergence rate is heterogeneous across iterations. Finally, we note that the rate is always guaranteed to be less than one because of Remark 5.4.
@@
Theorem 18 differentiates from [Banerjee et al., 2023, Theorem 5.3] in that (i) it is deterministic, i.e., it is not stated as holding with high probability; (ii) does not have all the weights of the network defined over a particular neighborhood (with the exception of the output linear layer). Indeed, these differences also hold for all the previously derived results in Section 4 and Section 5. ∎
@@ 6 Generalization guarantees for WeightNorm
We establish generalization bounds for WeightNorm networks. Our bound is based on a uniform convergence argument by bounding the Rademacher complexity of functions of the form (1) as long as the last layer vector $\mathbf{v}$ stays within a ball of radius $\rho_{1}$–the same condition we used in our optimization analysis. The core of the analysis relies on the use of a contraction-like property across the network layers similar to [Golowich et al., 2018] and uses WeightNorm and the network’s scaling to avoid exponential dependence on the depth and avoid any dependence on the width. All proofs are found in Section D of the appendix.
@@
[…]
@@
[…]
@@ 7 Experimental results
We show empirical results to support our theory. In our experiments, the WeightNorm networks are fully connected with two hidden layers of equal width $m$, $\operatorname{tanh}$ activation function, and linear output layer. We do empirical evaluations on CIFAR-10 [Krizhevsky, 2009] and MNIST [Deng, 2012]. Because of computational efficiency, we apply mini-batch stochastic gradient descent (SGD) with batch size 512 to optimize the WeightNorm networks under mean squared loss. In our experiments we are only concerned about the training of WeightNorm networks from an optimization perspective–the fact that the training error minimizes across epochs is enough for our purposes. Therefore, using a squared loss with a linear output in our experiments suffices–as opposed to cross-entropy loss with a softmax output layer. Our experiments were conducted on a computing cluster with AMD EPYC 7713 64-Core Processor and NVIDIA A100 Tensor Core GPU.
In our theory, the convergence rate as in Theorem 18 is directly proportional to the RSC parameter $\alpha_{\theta_{t}}$, and it benefits from a larger positive $\alpha_{\theta_{t}}$. According to equation (13), $\alpha_{\theta}$ becomes larger under two conditions: as the ratio $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$ increases and/or the minimum weight vector norm ${\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i,t}\right\rVert_{2}}$ increases. Then, the relevant question is: Can we provide empirical evidence of these two conditions benefiting convergence during training?
@@
Next, we focus on the effect of the minimum weight vector norm on the RSC parameter. According to our theory, if the two networks in Figure 1 have the same value for the ratio $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$, the network with the larger minimum weight vector norm will have a larger RSC parameter (see equation (13)), and so we would expect such network to have a faster convergence. Indeed, we observe this in our experiment: when the red solid curve crosses the dashed red curve (around epochs 70 and 115) in Figure 1(b), the two networks have the same value for $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$. At this crossing, Figure 1(b) shows that the network with larger width–and so with larger minimum weight vector norm–converges faster afterwards (i.e., the loss ratio ${\cal L}(\theta_{t+1})/{\cal L}(\theta_{t})$ for the green dashed curve has smaller values than the green solid curve).
@@ A.2 Upper bounds of the $L_{2}$-norms of $\alpha^{(l)}$
[…]
The $l=1$ case follows in a similar manner:
$\left\|\frac{\partial\alpha^{(1)}}{\partial\mathbf{w}^{(1)}}\right\|_{2}^{2}\leq\frac{1}{d}\|\alpha^{(0)}\|_{2}^{2}=\frac{1}{d}\left\lVert\mathbf{x}\right\rVert_{2}^{2}=1~{}$
which satisfies the form for $l=1$.
@@ Appendix B Upper bounds on gradients of the predictor and empirical loss for WeightNorm
[…]
where where (a) follows with probability $1-\frac{2}{m}(L+1)$ using a union bound from all the previously used results. We further note that,
@@ C.1 Restricted Strong Convexity and Smoothness
$\overset{(c)}{\geq}\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\frac{1}{n}\sum_{i=1}^{n}\bar{\ell}^{\prime}_{i}\frac{\partial f(\theta;\mathbf{x}_{i})}{\partial\theta}\right\rangle^{2}-\frac{4}{n}\sum_{i=1}^{n}\varrho\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\|\tilde{\theta}-\theta\|_{2}\|\theta^{\prime}-\theta\|_{2}^{2}$
$=\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\right\rangle^{2}-\frac{4}{n}\sum_{i=1}^{n}\varrho\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\|\tilde{\theta}-\theta\|_{2}\|\theta^{\prime}-\theta\|_{2}^{2}$
$\overset{(d)}{\geq}\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\right\rangle^{2}-4\varrho\|\theta^{\prime}-\theta\|_{2}^{3}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}$
$\overset{(e)}{\geq}\frac{\kappa^{2}}{2{\cal L}(\theta)}\left\|\nabla_{\theta}{\cal L}(\theta)\right\|_{2}^{2}\|\theta^{\prime}-\theta\|_{2}^{2}-4\varrho\|\theta^{\prime}-\theta\|_{2}^{3}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}$
$=\left(\frac{\kappa^{2}}{2{\cal L}(\theta)}\left\|\nabla_{\theta}{\cal L}(\theta)\right\|_{2}^{2}-4\varrho\|\theta^{\prime}-\theta\|_{2}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\right)\|\theta^{\prime}-\theta\|_{2}^{2}~{},$
@@
where (a) follows from the Cauchy-Schwartz inequality and the submultiplicative property of the induced norm; (b) follows by Cauchy-Schwartz inequality again; and (c) from $\frac{1}{n}\|[\ell^{\prime}_{i}]_{i}\|_{2}^{2}=\frac{4}{n}\sum_{i=1}^{n}(\tilde{y}_{i}-y_{i})^{2}=4{\cal L}(\theta)$.
@@
$\leq(4\varrho_{\theta}\rho_{2}+2\sqrt{\varphi}){\cal O}\left(\frac{(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\left\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\right\}}\right)$
$\leq{\cal O}\left((\varrho_{\theta}\rho_{2}+\sqrt{\varphi})\frac{(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\left\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\right\}}\right)$
@@
×$\left.\qquad\times\frac{\max\{(1+\rho_{1}),\rho_{2}\}(1+\rho_{1})^{2}L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})(1+L|\phi(0)|\sqrt{m})}{\min\left\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\right\}}\right)$
$\leq{\cal O}\left(\left(1+\frac{\sqrt{L}}{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}}}\right)\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})(1+L|\phi(0)|\sqrt{m})}{\min\left\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\right\}}\right)$
@@
By the second order Taylor expansion about $\bar{\theta}\in\mathbb{R}^{p}$ with $\bar{\mathbf{v}}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$, we have ${\cal L}(\theta^{\prime})={\cal L}(\bar{\theta})+\langle\theta^{\prime}-\bar{\theta},\nabla_{\theta}{\cal L}(\bar{\theta})\rangle+\frac{1}{2}(\theta^{\prime}-\bar{\theta})^{\top}\frac{\partial^{2}{\cal L}(\tilde{\theta})}{\partial\theta^{2}}(\theta^{\prime}-\bar{\theta})$, where $\tilde{\theta}=\xi\theta^{\prime}+(1-\xi)\bar{\theta}$ for some $\xi\in[0,1]$. Similarly to what was proved in Theorem 5.1, $\tilde{\mathbf{v}}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$.

### v2

@@ 5 Optimization guarantees for WeightNorm
AB-PCV-LZ-MB:22] Let $B_{t}:=Q^{\theta_{t}}_{\kappa}\cap B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})\cap\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}$, $\theta^{*}\in\operatorname{arginf}_{\theta\in\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}}{\cal L}(\theta)$, $\overline{\theta}_{t}\in\operatorname{arginf}_{\theta\in B_{t}}{\cal L}(\theta)$, and $\gamma_{t}:=\frac{{\cal L}(\overline{\theta}_{t})-{\cal L}(\theta^{*})}{{\cal L}(\theta_{t})-{\cal L}(\theta^{*})}$. Let $\alpha_{\theta_{t},\overline{\theta}_{t}},\beta_{\theta_{t}}$ be as in Lemmas 5.1 and 5.2 respectively with $\beta_{\theta_{t}}\equiv\beta_{\theta_{t},\theta_{t+1}}=\beta_{\theta_{t},\theta_{t}-\eta_{t}\nabla_{\theta}{\cal L}(\theta_{t})}$. Consider Assumptions 12, and 3, and the RSC property $\alpha_{\theta_{t},\overline{\theta}_{t}}>0$. Then, we have $\gamma_{t}\in[0,1)$; and if ${\cal L}(\theta_{t})\neq{\cal L}(\theta^{*})$, then for gradient descent with step size $\eta_{t}=\frac{1}{\beta_{\theta_{t}}}$, we have $\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}\in(0,1]$ and
${\cal L}(\theta_{t+1})-{\cal L}(\theta^{*})\leq\left(1-\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}(1-\gamma_{t})\right)({\cal L}(\theta_{t})-{\cal L}(\theta^{*}))~{}.$ (18)
The result in (18) implies training convergence as the number of iterations increase. Since the rate at which the loss decreases is time- and weight-dependent in (18), it is difficult to establish the number of steps needed for convergence to a specific loss value; however, this shows that our framework may cover cases where the convergence rate is heterogeneous across iterations. Finally, we note that the rate is always guaranteed to be less than one.
@@
Theorem 18 differentiates from [Banerjee et al., 2023, Theorem 5.3] in that (i) it is deterministic, i.e., it is not stated as holding with high probability; (ii) does not have all the weights of the network defined over a particular neighborhood (with the exception of the output linear layer). Indeed, these differences also hold for all the previously derived results in Section 4 and Section 5. We summarize the comparison between our results and those for networks without WeightNorm by Banerjee et al. [2023] in Table 1 from Section D of the appendix. ∎
Remark 5.6 (About the radius $\rho_{2}$ in Assumption 3).
We prove in Proposition C.2 of the appendix that, for the case of $\phi(0)=0$, $\left\lVert\theta_{t+1}-\theta_{t}\right\rVert_{2}\leq\eta_{t}\cdot{\cal O}\left((1+\rho_{1})^{2}\left(1+\frac{\sqrt{L}}{\sqrt{m}\left\lVert\bar{W}_{t}\right\rVert_{2}}\right)\right)$. Note that $\rho_{2}$ as in Assumption 3 could in principle depend on $t$, and thus we could set $\rho_{2}$ to be equal to the upper bound just shown for $\left\lVert\theta_{t+1}-\theta_{t}\right\rVert_{2}$. Remarkably, $\eta_{t}$ as chosen in Theorem 18 can be upper bounded by one [^1] We can set $\beta_{\theta_{t}}$ so that $1\leq\beta_{\theta_{t}}$ from (17) and so $\eta_{t}=\frac{1}{\beta_{\theta_{t}}}\leq 1$. and if $\left\lVert\bar{W}_{t}\right\rVert_{2}$ has a uniform lower bound across $t$ (note that our experiments in Section 7 provide evidence for $\left\lVert\bar{W}_{t}\right\rVert_{2}\geq\left\lVert\bar{W}_{0}\right\rVert_{2}$), then the upper bound of $\left\lVert\theta_{t+1}-\theta_{t}\right\rVert_{2}$ becomes a constant independent of $t$, and so does $\rho_{2}$ to satisfy (A3.2) in Assumption 3. ∎
Remark 5.7 (About gradient descent in Assumption 3).
Regarding (A3.1) in Assumption 3, we remark that we are only considering that the weight vector of the output layer is inside a neighborhood of its initialization value across iterations—no constraint exist on the weights from the hidden layers. Moreover, the radius $\rho_{1}$ of this neighborhood region can be set to be any arbitrary constant of order $\Theta(1)$ or even a polynomial of $L$; thus, $\rho_{1}$ is not assumed to be arbitrarily close to zero and the neighborhood is not restricted to be arbitrarily small. These arguments imply that our optimization guarantees are not restricted to be in the lazy training regime [Liu et al., 2020]. We also point out that (A3.1) is a necessary assumption so that the linear output of the neural network, which has no normalization, remains bounded. Finally, we remark that (A3.1) is not stronger than other existing works where assumptions of closeness around initialization points are needed even on the hidden weights; e.g., see [Liu et al., 2020, 2022, Banerjee et al., 2023]. ∎
@@ 6 Generalization guarantees for WeightNorm
We establish generalization bounds for WeightNorm networks. Our bound is based on a uniform convergence argument by bounding the Rademacher complexity of functions of the form (1) as long as the last layer vector $\mathbf{v}$ stays within a ball of radius $\rho_{1}$—the same condition we used in our optimization analysis. The core of the analysis relies on the use of a contraction-like property across the network layers similar to [Golowich et al., 2018] and uses WeightNorm and the network’s scaling to avoid exponential dependence on the depth and avoid any dependence on the width. All proofs are found in Section E of the appendix.
@@
Theorem 6.1, unlike our previous results in the paper, requires the smooth activation function to satisfy $\phi(0)=0$. As mentioned earlier in Remarks 4.2 and 5.3, the case when $\phi(0)=0$ improves our Hessian bounds and thus improves our optimization conditions by introducing a better dependence on the width $m$. We point out that commonly used activation functions such as the hyperbolic tangent function $\operatorname{tanh}$ and Gaussian Error Linear Unit ($\operatorname{GELU}$) satisfy both $\phi(0)=0$ and Assumption 1. ∎
@@
The generalizaton bound (19) does not explicitly depend on the network’s width, but explicitly depends on its depth. Setting $\rho_{1}=\Theta(1)$, a choice which does not negatively affect our optimization results, results in a generalization bound of ${\cal O}(\sqrt{\frac{L}{n}})$. Effectively, for deeper networks, we would need to increase the number of samples at least linearly on the depth to improve our generalization bound—in practice, the number of samples is usually much larger than the network’s depth. This is a benign dependence that WeightNorm allows compared to an exponential one found in networks without WeightNorm [Bartlett et al., 2017, Neyshabur et al., 2018, Golowich et al., 2018]. ∎
@@ 7 Experimental results
We show empirical results to support our theory. In our experiments, the WeightNorm networks are fully connected with two hidden layers of equal width $m$, $\operatorname{tanh}$ activation function, and linear output layer. We do empirical evaluations on CIFAR-10 [Krizhevsky, 2009] and MNIST [Deng, 2012]. Because of computational efficiency, we apply mini-batch stochastic gradient descent (SGD) with batch size 512 to optimize the WeightNorm networks under mean squared loss. In our experiments we are only concerned about the training of WeightNorm networks from an optimization perspective—the fact that the training error minimizes across epochs is enough for our purposes. Therefore, using a squared loss with a linear output in our experiments suffices—as opposed to cross-entropy loss with a softmax output layer. Our experiments were conducted on a computing cluster with AMD EPYC 7713 64-Core Processor and NVIDIA A100 Tensor Core GPU.
In our theory, the convergence rate as in Theorem 18 is directly proportional to the RSC parameter $\alpha_{\theta_{t},\overline{\theta}_{t}}$, and it benefits from a larger positive $\alpha_{\theta_{t},\overline{\theta}_{t}}$. According to equation (13), $\alpha_{\theta_{t},\overline{\theta}_{t}}$ becomes larger under two conditions: as the ratio $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$ increases and/or the minimum weight vector norm ${\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i,t}\right\rVert_{2}}$ increases. Then, the relevant question is: Can we provide empirical evidence of these two conditions benefiting convergence during training?
@@
Next, we focus on the effect of the minimum weight vector norm on the RSC parameter. According to our theory, if the two networks in Figure 1 have the same value for the ratio $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$, the network with the larger minimum weight vector norm will have a larger RSC parameter (see equation (13)), and so we would expect such network to have a faster convergence. Indeed, we observe this in our experiment: when the red solid curve crosses the dashed red curve (around epochs 70 and 115) in Figure 1(b), the two networks have the same value for $\|\nabla{\cal L}(\theta_{t})\|_{2}^{2}/{\cal L}(\theta_{t})$. At this crossing, Figure 1(b) shows that the network with larger width—and so with larger minimum weight vector norm—converges faster afterwards (i.e., the loss ratio ${\cal L}(\theta_{t+1})/{\cal L}(\theta_{t})$ for the green dashed curve has smaller values than the green solid curve).
@@ 8 Conclusion
Acknowledgments
Part of the work was done when Pedro Cisneros-Velarde was affiliated with the University of Illinois Urbana-Champaign and was concluded during his current affiliation with VMware Research. The work was supported in part by the National Science Foundation (NSF) through awards IIS 21-31335, OAC 21-30835, DBI 20-21898, as well as a C3.ai research award. Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, NIFA award 2020-67021-32799, the Alfred P. Sloan Foundation, and Google Inc.
@@ Appendix B Upper bounds on gradients of the predictor and empirical loss for WeightNorm
We further note that,
@@ C.1 Restricted Strong Convexity and Smoothness
$\overset{(c)}{\geq}\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\frac{1}{n}\sum_{i=1}^{n}\bar{\ell}^{\prime}_{i}\frac{\partial f(\theta;\mathbf{x}_{i})}{\partial\theta}\right\rangle^{2}-\frac{4}{n}\sum_{i=1}^{n}\varrho_{\theta}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\|\tilde{\theta}-\theta\|_{2}\|\theta^{\prime}-\theta\|_{2}^{2}$
$=\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\right\rangle^{2}-\frac{4}{n}\sum_{i=1}^{n}\varrho_{\theta}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\|\tilde{\theta}-\theta\|_{2}\|\theta^{\prime}-\theta\|_{2}^{2}$
$\overset{(d)}{\geq}\frac{1}{2{\cal L}(\theta)}\left\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\right\rangle^{2}-4\varrho_{\theta}\|\theta^{\prime}-\theta\|_{2}^{3}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}$
$\overset{(e)}{\geq}\frac{\kappa^{2}}{2{\cal L}(\theta)}\left\|\nabla_{\theta}{\cal L}(\theta)\right\|_{2}^{2}\|\theta^{\prime}-\theta\|_{2}^{2}-4\varrho_{\theta}\|\theta^{\prime}-\theta\|_{2}^{3}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}$
$=\left(\frac{\kappa^{2}}{2{\cal L}(\theta)}\left\|\nabla_{\theta}{\cal L}(\theta)\right\|_{2}^{2}-4\varrho_{\theta}\|\theta^{\prime}-\theta\|_{2}\cdot\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\bar{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}\right)\|\theta^{\prime}-\theta\|_{2}^{2}~{},$
@@
where (a) follows from the Cauchy-Schwartz inequality and the submultiplicative property of the induced norm; (b) follows by Cauchy-Schwartz inequality again; and (c) from $\frac{1}{n}\|[\ell^{\prime}_{i}]_{i}\|_{2}^{2}=\frac{4}{n}\sum_{i=1}^{n}(\tilde{y}_{i}-y_{i})^{2}=4{\cal L}(\tilde{\theta})$.
@@
$\leq(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}){\cal O}\left(\frac{4\varrho_{\theta}\rho_{2}}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2}^{2}\right\}}\right.$
$\left.\qquad+\frac{2\sqrt{\varphi}}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}\right\}}\right)$
$\leq(4\varrho_{\theta}\rho_{2}+2\sqrt{\varphi}){\cal O}\left(\frac{(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2}^{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\right\}}\right)$
$\leq{\cal O}\left((\varrho_{\theta}\rho_{2}+\sqrt{\varphi})\frac{(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2}^{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\right\}}\right)$
@@
×$\left.\qquad\times\frac{\max\{(1+\rho_{1}),\rho_{2}\}(1+\rho_{1})^{2}L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})(1+L|\phi(0)|\sqrt{m})}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2}^{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\right\}}\right)$
$\leq{\cal O}\left(\left(1+\frac{\sqrt{L}}{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}}}\right)\right.$
×$\left.\qquad\times\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})(1+L|\phi(0)|\sqrt{m})}{\min\left\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\bar{W}_{i}^{(l)}\right\rVert_{2}^{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\right\}}\right)$
@@
By the second order Taylor expansion about $\theta\in\mathbb{R}^{p}$ with $\bar{\mathbf{v}}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$, we have ${\cal L}(\theta^{\prime})={\cal L}(\theta)+\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\rangle+\frac{1}{2}(\theta^{\prime}-\theta)^{\top}\frac{\partial^{2}{\cal L}(\tilde{\theta})}{\partial\theta^{2}}(\theta^{\prime}-\theta)$, where $\tilde{\theta}=\xi\theta^{\prime}+(1-\xi)\theta$ for some $\xi\in[0,1]$. Similarly to what was proved in Theorem 5.1, $\tilde{\mathbf{v}}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$.